OpenStudy (anonymous):

Can someone please walk me through dividing these fractions? The answers have to be improper fractions. 4/5 / 1/6= 5/6 / 1/6 = 5/6 / 1/3= 7/8 / 1/10= 1/10 / 2/5= 1/2 / 3/10=

OpenStudy (anonymous):

\[\frac{4}{5}\div \frac{1}{6}=\frac{4}{5}\times \frac{6}{1}=\frac{4\times 6}{5\times 1}=\frac{24}{5}\] others are similar

OpenStudy (anonymous):

The way that I learned how to divide fractions is to just flip-an-multiply. instead of trying to dived each number flip the second fraction and multiply across.